1) In a saturated solution of Zn(OH)2 at 25°C, the value of [OH-] is 2.0 x 10-4 M. What is the value of the solubility-product constant, Ksp, for Zn(OH)2 at 25°C ? Explain

A)4.0 x 10, -18

B)8.0 x 10, -12

C)4.0 x 10-12

D)2.0 x 10-4


2) CaF2(s)⇄Ca2+(aq)+2F(aq)          Ksp=4.0×10−10

The concentration of F(aq) in drinking water that is considered to be ideal for promoting dental health is 4.0×10−5M. Based on the information above, the maximum concentration of Ca2+(aq) that can be present in drinking water without lowering the concentration of F(aq) below the ideal level is closest to

  • 0.025 M
  • 1.6 x 10-15 M
  • 1.6 x 10-6 M
  • 0.25 M


*EXPLAIN*


1
Expert's answer
2021-03-31T08:05:16-0400

Solution:

(1):

Zn(OH)2(s) ⇄ Zn2+(aq) + 2OH(aq)

___x_______ ___x__________2x_____

Hence,

[OH] = 2x = 2.0×10-4 M

[Zn2+] = x = [OH] / 2 = (2.0×10-4 M) / 2 = 1.0×10-4 M


The Ksp expression for zinc hydroxide is:

Ksp = [Zn2+][OH]2

Ksp = (1.0×10-4) × (2.0×10-4)2 = 4×10-12

Ksp = 4×10-12


NB!: Ksp for Zn(OH)2 is 5.0×10-17 (Table value. Maybe there is a mistake in the problem statement).


Answer (1): C) 4×10-12


(2):

CaF2(s) ⇄ Ca2+(aq) + 2F(aq)

The Ksp expression for calcium fluoride is:

Ksp = [Ca2+][F]2


Ksp= 4.0×10−10

[F] = 2x = 4.0×10−5 M


Hence,

[Ca2+]max = Ksp / [F]2

[Ca2+]max = (4.0×10−10) / (4.0×10−5)2 = 0.25

[Ca2+]max = 0.25 M


Answer (2): [Ca2+]max = 0.25 M.